# SFVInt: Simple, Fast and Generic Variable-Length Integer Decoding using Bit Manipulation Instructions

### v1

@@ 4. Performance Evaluation
Figure 4. Workload 1 (W1): uniform distribution - generated by folly::Random::secureRand32() or secureRand64().
@@
In our evaluation of the BMI2-enhanced varint decoding technique, we leveraged various AWS EC2 instances with diverse Intel and AMD CPU architectures (refer to Table 1). We created datasets encoded as varints, following uniform and skewed distributions. The uniform distribution served as a stable benchmark, while the skewed distribution, biased towards smaller integers (1-byte or 2-bytes LEBs), mirrored real-world varint patterns.
We employed four distinct workloads for evaluation: W1 featured uniformly distributed 32-bit integers as a benchmark; W2 used real-world LEB lengths from the WebAssembly build suite (web, [n. d.]), providing practical insights; and Workloads W3 and W4 derived from our proprietary systems’ data distributions, offering a realistic view of their application. Notably, W2 to W4 exhibited a skew towards 1-byte integers, aligning with typical data trends observed in practical scenarios. The specific distributions for these workloads are detailed in the figure captions of Figures 4 to 7.
@@ 4.1. Intel CPU Microarchitectures
In Figures 4-7, the left panel showcases the varint decoding performance across six EC2 instance types featuring Intel CPUs, detailed in Table 1. Each figure within the panel includes benchmarks for each EC2 instance type, with the y-axis representing the time required per iteration (in milliseconds), where each iteration involves decoding one million integers generated according to a specific distribution. Here, SFVInt consistently excels, surpassing all other systems in decoding speed across every instance type and under various distributions. For example, in W1 (Figure 4), SFVInt demonstrates a significant performance edge, nearly doubling the speed of Protobuf. Similarly, in W4 (Figure 7), SFVInt outperforms all other systems by nearly a factor of two.
@@ 4.2. AMD CPU Microarchitectures
In our analysis detailed in Figures 4-7, we observe that SFVInt demonstrates up to 40% faster varint decoding on 3rd generation AMD EPYC processors (m6a and c6a), in contrast to the slower performance on 2nd generation CPUs (m5a and c5a). This performance disparity is consistent with the known emulation overhead of PEXT/PDEP instructions on AMD platforms, which can significantly lag behind Intel’s native execution (Kullberg, 2017; red, 2017). Studies indicate that on AMD hardware, these instructions’ latency can vary widely, from 18 to approximately 300 cycles (InstLatX64, 2019). The improved results with newer AMD generations suggest advancements in BMI2 instruction handling. For enhanced consistency across CPU architectures, we can incorporate a dynamic selection mechanism between BMI2-accelerated and standard decoding or employs ZP7 (Wegner, 2019), a branchless alternative for AMD processors.
@@ 6. Conclusions
In this paper, we introduced SFVInt, a simple, fast and generic varint decoding approach leveraging BMI2 instructions, significantly enhancing the performance of varint decoding. Through rigorous evaluation across diverse CPU architectures, SFVInt consistently demonstrated up to a 2x increase in decoding speed over traditional methods on modern processors. This performance boost is crucial for applications dealing with large-scale data processing.

### v3

@@ 4. Performance Evaluation
Figure 5. Workload 1 (W1): uniform distribution.
@@
In our comprehensive evaluation of the BMI2-enhanced varint decoding technique, we utilized a diverse set of AWS EC2 instances equipped with a range of Intel and AMD CPU architectures (refer to Table 1). This heterogeneous testing environment allowed us to assess the performance and versatility of our approach across different processor generations and microarchitectures. To thoroughly evaluate the efficiency of our technique, we generated datasets encoded as varints following two distinct distributions: uniform and skewed. The uniform distribution dataset served as a stable benchmark, providing a balanced representation of integer values across the entire range. This dataset enabled us to assess the performance of our approach under a consistent and predictable workload. On the other hand, the skewed distribution dataset was specifically designed to mirror real-world varint patterns commonly encountered in various applications and data formats. This dataset exhibited a bias towards smaller integers, with a higher frequency of 1-byte and 2-byte LEBs. By incorporating a skewed distribution, we aimed to evaluate the performance of our technique in scenarios that closely resemble practical use cases.
We employed four distinct workloads for evaluation: W1 featured uniformly distributed 32-bit integers as a benchmark; W2 used real-world LEB lengths from the WebAssembly build suite (web, [n. d.]), providing practical insights; and Workloads W3 and W4 derived from our proprietary systems’ data distributions, offering a realistic view of their application. Notably, W2 to W4 exhibited a skew towards 1-byte integers, aligning with typical data trends observed in practical scenarios. The specific distributions for these workloads are detailed in the figure captions of Figures 5 to 8.
@@ 4.1. Intel CPU Microarchitectures
In Figures 5-8, the left panel showcases the varint decoding performance across six EC2 instance types featuring Intel CPUs, detailed in Table 1. Each figure within the panel includes benchmarks for each EC2 instance type, with the y-axis representing the time required per iteration (in milliseconds), where each iteration involves decoding one million integers generated according to a specific distribution. Here, SFVInt consistently excels, surpassing all other systems in decoding speed across every instance type and under various distributions. For example, in W1 (Figure 5), SFVInt demonstrates a significant performance edge, nearly doubling the speed of Protobuf. Similarly, in W4 (Figure 8), SFVInt outperforms all other systems by nearly a factor of two.
@@ 4.2. AMD CPU Microarchitectures
In our analysis detailed in Figures 5-8, we observe that SFVInt demonstrates up to 40% faster varint decoding on 3rd generation AMD EPYC processors (m6a and c6a), in contrast to the slower performance on 2nd generation CPUs (m5a and c5a). This performance disparity is consistent with the known emulation overhead of PEXT/PDEP instructions on AMD platforms, which can significantly lag behind Intel’s native execution (Kullberg, 2017; red, 2017). Studies indicate that on AMD hardware, these instructions’ latency can vary widely, from 18 to approximately 300 cycles (InstLatX64, 2019). The improved results with newer AMD generations suggest advancements in BMI2 instruction handling. For enhanced consistency across CPU architectures, we can incorporate a dynamic selection mechanism between BMI2-accelerated and standard decoding or employs ZP7 (Wegner, 2019), a branchless alternative for AMD processors.
@@ 5. Related Works
During our investigation of next-generation columnar store for machine learning workloads (Liao et al., 2024), we found discussions on Google Protobuf regarding the use of BMI2 for accelerating varint decoding (Zhou, 2024; Palade, 2024). The Protobuf team expressed concerns about the performance of BMI2 on AMD processors. Motivated by these discussions, we developed SFVInt and conducted a comprehensive analysis of the effectiveness of BMI2 on both Intel and AMD platforms, addressing the performance considerations raised by the community.
@@ 6. Conclusions
In this paper, we presented SFVInt, a simple, fast, and generic approach for decoding variable-length integers using BMI2 instructions. Through extensive evaluation across diverse CPU architectures, SFVInt consistently demonstrated up to a 2x increase in decoding speed compared to traditional methods, addressing a critical bottleneck in data-intensive applications. As data continues to grow in volume and complexity, SFVInt represents a step forward in offering a high-performance solution for varint decoding that can greatly benefit a wide range of applications.